# An Accelerated Hybrid Low-Rank Augmented Lagrangian Method for Large-Scale Semidefinite Programming on GPU

### v1

@@ Abstract
This paper introduces cuHALLaR, a GPU-accelerated implementation of the HALLaR method proposed in [47] for solving large-scale semidefinite programming (SDP) problems. We demonstrate how our Julia-based implementation efficiently uses GPU parallelism through optimization of simple, but key, operations, including linear maps, adjoints, and gradient evaluations. Extensive numerical experiments across three problem classes—maximum stable set, matrix completion, and phase retrieval show significant performance improvements over both CPU implementations and existing GPU-based solvers. For the largest instances, cuHALLaR achieves speedups of 30-140x on matrix completion problems, up to 135x on maximum stable set problems for Hamming graphs with 8.4 million vertices, and 15-47x on phase retrieval problems with dimensions up to 3.2 million. Our approach efficiently handles massive problems with dimensions up to $(n,m)\approx(8\times 10^{6},3\times 10^{8})$ at $10^{-5}$ relative precision, solving matrix completion instances with over 8 million rows and columns in just 142 seconds. These results establish cuHALLaR as a very promising GPU-based method for solving large-scale semidefinite programs.
@@ 1 Introduction
Let $\mathbb{S}^{n}$ be the set of $n\times n$ symmetric matrices. The notation $A\succeq B$ means that $A-B$ is positive semidefinite. We are interested in solving the primal-dual pair of semidefinite programs (SDPs)
$P_{*}:=\quad\min_{X}\quad\{C\bullet X\quad:\quad\mathcal{A}(X)=b,\quad X\in\Delta^{n}\}$ (P)
@@
$D_{*}:=\quad\max_{p\in\mathbb{R}^{m},\theta\in\mathbb{R}}\quad\{-b^{\top}p-\theta\quad:\quad S\vcentcolon=C+\mathcal{A}^{*}(p)+\theta I\succeq 0,\quad\theta\geq 0\}$ (D)
where $b\in\mathbb{R}^{m}$, $C\in\mathbb{S}^{n}$, $\mathcal{A}:\mathbb{S}^{n}\to\mathbb{R}^{m}$ is a linear map, $\mathcal{A}^{*}:\mathbb{R}^{m}\to\mathbb{S}^{n}$ is its adjoint, and $\Delta^{n}$ is the spectraplex
$\Delta^{n}\vcentcolon=\{X\in\mathbb{S}^{n}:\mathrm{Tr}(X)\leq 1,X\succeq 0\}.$ (1)
@@ 1.1 Related literature
We divide our discussion into four parts: first, the emerging literature on utilizing GPUs for mathematical optimization, second, the challenges in solving large SDPs, third, recent techniques in solving large SDPs using low-rank approaches, and fourth, a comparison between cuHALLaR, a recent GPU SDP solver cuLoRads, and the CPU version of HALLaR.
@@
GPUs offer massive parallelism that accelerates mathematical programming through parallel linear algebra operations. Recent GPU-based solvers for linear programming and quadratic programming [41, 43, 42] have achieved performance which outperforms traditional CPU-based methods. First-order methods map well to GPU architectures, unlike simplex or interior-point methods which rely on difficult-to-parallelize factorizations. For linear programming, cuPDLP.jl incorporates GPU-accelerated sparse matrix operations to match commercial solver performance. Quadratic programming solvers like cuOSQP [57] and CuClarabel [12] similarly exploit GPU parallelism for significant speedups. Most recently, [39] proposed a new GPU conic programming solver that utilizes the primal-dual hybrid gradient (PDHG) method for solving several classes of cone optimization problems, including semidefinite programs.
@@
The low-rank approach proposed by Burer and Monteiro [9, 10] converts SDP (P) into a nonconvex quadratic program with quadratic constraints (QCQP) through the transformation $X=UU^{T}$, where $U\in\mathbb{R}^{n\times r}$ and $r\ll n$. The main advantage of this approach is that the QCQP has $n\times r$ variables, while (P) has $n(n+1)/2$ variables. Also, the PSD constraint $X\succeq 0$ has been eliminated through the low-rank transformation so methods do not need to perform eigendecompositions of dense matrices to approximately solve the QCQP. Burer and Monteiro proposed an efficient first-order limited-memory BFGS augmented Lagrangian (AL) method, which only relied on matrix-vector multiplications and storage of $n\times r$ matrices, to find an approximate local minimizer of the QCQP. Recent landscape results have shown that if the factorization parameter $r$ is taken large enough, then finding local minimizers of the QCQP is actually equivalent to finding global minimizers of the SDP (P) [14, 15, 7, 8, 4, 55, 67]. For a more extensive list of papers which characterize the optimization landscape of nonconvex formulations obtained through Burer-Monteiro factorizations, see [50, 22, 44, 40, 49, 18, 45, 37]. For recent advances and developments in accelerating low-rank first-order methods for solving SDPs, we refer the reader to the extensive literature available in [23, 29, 30, 47, 62, 63, 64, 69, 70, 77, 68, 76, 78].
@@
We now give a brief comparison between HALLaR, which was first developed in [47] and the LoRADS method developed in [29]. HALLaR utilizes an inexact AL method to solve (P). It solves each of its AL subproblems over the spectraplex $\operatorname*{arg\,min}_{X\in\Delta^{n}}\quad C\bullet X+p^{\top}(\mathcal{A}(X)-b)+\frac{\beta}{2}\|\mathcal{A}X-b\|^{2}$ by first restricting them to the space of matrices of rank at most $r$ through the transformation $X=UU^{T}$ and then applying a nonconvex solver to the reformulation. It uses a low-rank Frank-Wolfe step to escape from a possible spurious stationary point obtained by the nonconvex solver, a step which also potentially increases the rank of the current iterate.
@@
LoRADS, on the other hand, adopts a two-stage approach. In their first stage, they utilize the low-rank limited-memory BFGS AL method of Burer and Monteiro as a warm-start strategy for their second stage. The second stage of LoRADS also utilizes an inexact AL method applied to (P) but the major difference being the way the AL subproblems are reformulated. In contrast to the $X=UU^{T}$ change of variables used by HALLaR, LoRADS reformulates them by replacing $X$ by $UV^{T}$ and adding the constraint $U=V$. It then approximately solves these nonconvex reformulations of the AL subproblems by an alternating minimization penalty-type approach which penalizes $U=V$ constraint and alternately minimizes with respect to the blocks $U$ and $V$ using the conjugate gradient method. Similar to HALLaR, the inexact stationary solutions of the nonconvex AL reformulations may not be near global solutions of the AL subproblems as they are not necessarily equivalent. In contrast to HALLaR, LoRADS uses a heuristic that increases the number of columns of $U$ and $V$ to attempt to escape from the these spurious solutions.
@@
Like cuLoRads, which is a GPU accelerated version of LoRads, the cuHALLaR method proposed in this paper is a GPU accelerated version of HALLaR. It is displayed on several important classes of SDPs that cuHALLaR is 2 to 25 times faster than cuLoRads on many large instances. cuHALLaR can efficiently solve massive problems to $10^{-5}$ relative precision in just a few minutes. For example, cuHALLaR is able to solve a matrix completion SDP instance where the size of the matrix variable is 8 million and the number of constraints is approximately 300 million in just 142 seconds.
@@ 1.2 Notation
[…]
@@ 2 Overview of HALLaR
As mentioned in the Introduction, HALLaR is an inexact AL method that approximately solves the pair of SDPs (P) and (D). Given a tolerance pair $(\epsilon_{p},\epsilon_{pd})\in\mathbb{R}^{2}_{++}$, HALLaR aims to find a $(\epsilon_{p},\epsilon_{pd})$-optimal solution of (P) and (D), i.e., a triple $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$ that satisfies
$\|\mathcal{A}\bar{X}-b\|\leq\epsilon_{p},\quad|C\bullet\bar{X}-(b^{T}\bar{p}-\bar{\theta})|\leq\epsilon_{pd},\quad C+\mathcal{A}^{*}p+\bar{\theta}I\succeq 0,\quad\bar{\theta}\geq 0.$ (3)
Being an inexact AL method, HALLaR generates sequences $\{X_{t}\}$ and $\{p_{t}\}$ according to the following updates
$X_{t}\quad$ $\approx\quad\operatorname*{arg\,min}_{X}\quad\{\mathcal{L}_{\beta_{t}}(X;p_{t-1})\;:\;X\in\Delta^{n}\},$ (4a)
@@
is the AL function. The key part of HALLaR is an efficient method, called the hybrid low-rank method (HLR), for finding a near-optimal solution $X_{t}$ of the AL subproblem (4a). HLR never forms $X_{t}$ but outputs a low-rank factor $U_{t}\in B_{1}^{s_{t}}$ such that $X_{t}=U_{t}U_{t}^{T}$ where $B_{1}^{s_{t}}$ is as in (2) and hopefully $s_{t}\ll n$. The next paragraph briefly describes HLR in more detail for solving (4a).
Given a pair $(s,\tilde{Y})\in\mathbb{Z}_{+}\times B_{1}^{s}$, which is initially set to $(s,\tilde{Y})=(s_{t-1},U_{t-1})$, a general iteration of HLR performs the following steps: i) it computes a suitable stationary point $Y$ of the nonconvex reformulation of (4a) obtained through the change of variable $X=UU^{T}$, namely:
$\min_{U}\quad\left\{g_{t}(U)=\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad U\in B_{1}^{s}\right\};$ ($\mathrm{L}_{r}$)
(ii) checks whether $X=YY^{\top}$, is a reasonable approximate solution of (4a); if so it stops and sets $(s_{t},U_{t})=(s,Y)$; and; iii) else, it performs a Frank-Wolfe step in the $X$-space (but implemented in the $U$-space) to obtain a new pair $(s,\tilde{Y})$.
We now give more details about the above steps. First, we consider step i). This step is implemented with the aid of a generic nonlinear solver which, started from $\tilde{Y}$, computes a $(\rho;\tilde{Y})$-stationary solution of ($\mathrm{L}_{r}$), i.e., a triple $(Y;R,\rho)$ such that
$R\in\nabla g_{t}(Y)+N_{B^{s}_{1}}(Y),\quad\|R\|_{F}\leq\rho,\quad g_{t}(Y)\leq g_{t}(\tilde{Y}).$ (6)
@@
$R\in 2[C+\mathcal{A}^{*}(p_{t-1}+\beta_{t}(\mathcal{A}(YY^{T})-b))]Y+N_{B^{s}_{1}}(Y),\quad\|R\|_{F}\leq\rho.$
@@
Algorithm 1 HALLaR. Output: $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$, an $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})$-solution of the pair of SDPs (P) and (D).
1:Let initial points $(U_{0},p_{0})\in B^{s_{0}}_{1}\times\mathbb{R}^{m}$, and tolerance pair $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})\in\mathbb{R}^{2}_{++}$ be given, and let $\left\{\epsilon_{t}\right\}_{t\geq 1}$ and $\left\{\beta_{t}\right\}_{t\geq 1}$ be a decreasing and increasing sequence of positive integers, respectively.
2:set $t\leftarrow 1$;
3:set $\tilde{Y}=U_{t-1}$, $s=s_{t-1}$, $Y=0_{n\times s}$, $G=0_{n\times n}$, and $\theta=\infty$;
[…]
5:     call a nonconvex solver with initial point $\tilde{Y}$, tolerance $\epsilon_{t}$, and function $\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})+\delta_{B_{1}^{s}}(U)$ to find a point $Y\in B^{s}_{1}$ that is an $\epsilon_{t}$-approximate stationary solution (according to the criterion in (6)) of
$\min_{U}\quad\left\{\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad\|U\|_{F}\leq 1,\quad U\in\mathbb{R}^{n\times s}\right\};$ (7)
6:     compute
$G:=\nabla(\mathcal{L}_{\beta_{t}}(YY^{T};p_{t-1}))\in\mathbb{S}_{n}$ (8)
@@
7:     if $(GY)\bullet Y+\theta\leq\epsilon_{t}$ then
8:         break while loop and go to step 11 below;
9:     end if
10:     compute
@@
and set
$(\tilde{Y},s)=\begin{cases}(y,1)&\text{ if $\alpha=1$}\\ \left(\left[\sqrt{1-\alpha}\,Y,\sqrt{\alpha}\,{y}\right],s+1\right)&\text{ otherwise};\end{cases}$ (11)
11:end while
12:set $(U_{t},\theta_{t},s_{t})=(Y,\theta,s)$;
13:set
$p_{t}=p_{t-1}+\beta(\mathcal{A}(U_{t}U_{t}^{\top})-b);$ (12)
14:if
$\|\mathcal{A}(U_{t}U_{t}^{\top})-b\|\leq\epsilon_{\mathrm{p}},\quad|CU_{t}\bullet U_{t}-(-b^{\top}p_{t}-\theta_{t})|\leq\epsilon_{pd}$ (13)
then return $(\bar{U},\bar{p}_{,}\bar{\theta})=(U_{t},p_{t},\theta_{t})$;
15:set $t=t+1$ and go to step 2.
Several remarks about each of the steps of HALLaR are now given. First, $t$ is the iteration count for HALLaR. Second, each iteration of HALLaR performs a HLR call and each call is performed in the while loop from steps 3 to 10. In line 2, the triple $(Y,G,\theta)$ is set as $(0_{n\times s},0_{n\times n},\infty)$, which means the HLR call in the while loop is always performed during each iteration of HALLaR. Third, after a HLR call is finished, HALLaR updates the Lagrange multiplier in line 12 and checks for termination in line 13. Lastly, the termination condition on this line does not check for near dual feasibility since it can be shown that for every $t\geq 1$,
@@
and hence that HALLaR always generates feasible dual iterates $(p_{t},\theta_{t},S_{t})$’s.
@@
We now comment on the steps performed within the while loop (lines 3 to 10) in light of the three steps of the HALLaR outline given just before its formal description. Step 4 is implementing step i) of the HALLaR outline. Steps 5 and 6 are implementing step ii). Finally, step 9 implements step iii).
We now give an interpretation of step 9 of HALLaR relative to the $X$-method. The stepsize $\alpha$ computed in (10) corresponds to solving
$\operatorname*{arg\,min}_{\alpha}\left\{\mathcal{L}_{\beta_{t}}\left(\alpha yy^{T}+(1-\alpha)YY^{T};p_{t-1}\right):\alpha\in[0,1]\right\}\}.$

### v2

# Authors are listed in alphabetical order

@@ Abstract
This paper presents an SDP solver, cuHALLaR, which is a GPU-accelerated implementation of the hybrid low-rank augmented Lagrangian method, HALLaR, proposed earlier by three of the authors. The proposed Julia-based implementation is designed to exploit massive parallelism by performing all core operator evaluations—linear maps, their adjoints, and gradients—entirely on the GPU. Extensive numerical experiments are conducted on three problem classes: matrix completion, maximum stable set, and phase retrieval. The results demonstrate substantial performance gains over both optimized CPU implementations and existing GPU-based solvers. On the largest instances, cuHALLaR achieves speedups of up to 165x, solving a matrix completion SDP with a matrix variable of size $2\times 10^{6}\times 2\times 10^{6}$ and over $2.6\times 10^{8}$ constraints to a relative precision of $10^{-5}$ in 53 seconds. Similarly, speedups of up to 135x are observed for maximum stable set problems on graphs with over one million vertices. These results establish cuHALLaR as a state-of-the-art solver for extremely large-scale SDPs.
@@ 1 Introduction
Let $\mathbb{S}^{n}$ be the set of $n\times n$ symmetric matrices. The notation $A\succeq B$ means that ABA-B is positive semidefinite. We are interested in solving the primal-dual pair of semidefinite programs (SDPs)
$P_{*}:=\quad\min_{X}\quad\{C\bullet X\quad:\quad\mathcal{A}(X)=b,\quad X\in\Delta^{n}_{\tau}\}$ (P)
@@
$D_{*}:=\quad\max_{p\in\mathbb{R}^{m},\theta\in\mathbb{R}_{+}}\quad\{-b^{\top}p-\tau^{2}\theta\quad:\quad S\vcentcolon=C+\mathcal{A}^{*}(p)+\theta I\succeq 0\}$ (D)
where $\bullet$ is the Frobenius inner product, $b\in\mathbb{R}^{m}$, $C\in\mathbb{S}^{n}$, and $\mathcal{A}:\mathbb{S}^{n}\to\mathbb{R}^{m}$ and $\mathcal{A}^{*}:\mathbb{R}^{m}\to\mathbb{S}^{n}$ are linear maps such that
$\mathcal{A}(X)=\begin{bmatrix}A_{1}\bullet X\\ A_{2}\bullet X\\ \vdots\\ A_{m}\bullet X\end{bmatrix},\quad\mathcal{A}^{*}(p)=\sum_{\ell=1}^{m}p_{\ell}A_{\ell},\quad\text{where $A_{\ell}\in\mathbb{S}^{n}$ for $\ell=1,\ldots m$}.$ (1)
We also define $\Delta^{n}_{\tau}$ to be the spectraplex
$\Delta^{n}_{\tau}\vcentcolon=\{X\in\mathbb{S}^{n}:\mathrm{Tr}(X)\leq\tau^{2},X\succeq 0\}.$ (2)
@@ 1.1 Related literature
We divide our discussion into four parts: first, the emerging literature on utilizing GPUs for mathematical optimization, second, the challenges in solving large SDPs, third, recent techniques in solving large SDPs using low-rank approaches, and fourth, a comparison between cuHALLaR, a recent GPU SDP solver cuLoRADS developed in [33], and the CPU version of HALLaR.
@@
GPUs offer massive parallelism that accelerates mathematical programming through parallel linear algebra operations. Recent GPU-based solvers for linear programming and quadratic programming [44, 46, 45] have achieved performance which outperforms traditional CPU-based methods. First-order methods map well to GPU architectures, unlike simplex or interior-point methods which rely on difficult-to-parallelize factorizations. For linear programming, cuPDLP.jl incorporates GPU-accelerated sparse matrix operations to match commercial solver performance. Quadratic programming solvers like cuOSQP [60] and CuClarabel [13] similarly exploit GPU parallelism for significant speedups.
Most recently, [42] proposed a new GPU primal-dual conic programming solver, namely, PDCS, which utilizes the primal-dual hybrid gradient (PDHG) method for solving several classes of cone optimization problems, including semidefinite programs; however, the authors of the paper remark that due to PDHG requiring Euclidean projections, their method for solving SDP problems is rather inhibited and slow. Thus, their code does not allow users to input SDP data, and so we cannot compare against PDCS.
@@
The low-rank approach proposed by Burer and Monteiro [10, 11] converts SDP (P) into a nonconvex quadratic program with quadratic constraints (QCQP) through the transformation $X=UU^{\top}$, where $U\in\mathbb{R}^{n\times r}$ and $r\ll n$. The main advantage of this approach is that the QCQP has $n\times r$ variables, while (P) has $n(n+1)/2$ variables. Also, the PSD constraint $X\succeq 0$ has been eliminated through the low-rank transformation so methods do not need to perform eigendecompositions of dense matrices to approximately solve the QCQP. Burer and Monteiro proposed an efficient first-order limited-memory BFGS augmented Lagrangian (AL) method, which only relied on matrix-vector multiplications and storage of $n\times r$ matrices, to find an approximate local minimizer of the QCQP. Recent landscape results have shown that if the factorization parameter $r$ is taken large enough, then finding local minimizers of the QCQP is actually equivalent to finding global minimizers of the SDP (P) [15, 16, 8, 9, 5, 58, 69]. For a more extensive list of papers which characterize the optimization landscape of nonconvex formulations obtained through Burer-Monteiro factorizations, see [53, 24, 47, 43, 52, 20, 48, 40]. For recent advances and developments in accelerating low-rank first-order methods for solving SDPs, we refer the reader to the extensive literature available in [25, 32, 33, 50, 64, 65, 66, 71, 72, 78, 70, 77, 79].
@@
We now give a brief comparison between HALLaR, which was first developed in [50] and the LoRADS method developed in [32]. HALLaR utilizes an inexact AL method to solve (P). It solves a sequence of AL subproblems of the form $\operatorname*{arg\,min}_{X\in\Delta^{n}_{\tau}}\ C\bullet X+p^{\top}(\mathcal{A}(X)-b)+\beta\|\mathcal{A}X-b\|^{2}/2$ by first restricting them to the space of matrices of the form $X=UU^{\top}$, where the number of columns $r$ of $U$ is significantly smaller than $n$, and then applying a nonconvex solver to the reformulation. It uses a low-rank Frank-Wolfe step to escape from a possible spurious stationary point obtained by the nonconvex solver, a step which generally adds an additional column to $U$.
@@
LoRADS, on the other hand, adopts a two-stage approach. In their first stage, they utilize the low-rank limited-memory BFGS AL method of Burer and Monteiro as a warm-start strategy for their second stage. The second stage of LoRADS also utilizes an inexact AL method applied to (P) but the major difference being the way the AL subproblems are reformulated. In contrast to the $X=UU^{\top}$ change of variables used by HALLaR, LoRADS reformulates them by replacing $X$ by $UV^{\top}$ and adding the constraint $U=V$. It then approximately solves these nonconvex reformulations of the AL subproblems by an alternating minimization penalty-type approach which penalizes $U=V$ constraint and alternately minimizes with respect to the blocks $U$ and $V$ using the conjugate gradient method. Similar to HALLaR, the inexact stationary solutions of the nonconvex AL reformulations may not be near global solutions of the AL subproblems as they are not necessarily equivalent. In contrast to HALLaR, LoRADS uses a heuristic that increases the number of columns of $U$ and $V$ to attempt to escape from the these spurious solutions.
@@
Like cuLoRads, which is a GPU accelerated version of LoRads developed in [33], cuHALLaR is also a GPU accelerated version of HALLaR. Computational results reported in this paper show that cuHALLaR is 2 to 25 times faster than cuLoRads on large-scale instances of some important SDP classes. cuHALLaR can efficiently solve massive problems to $10^{-5}$ relative precision in just a few minutes. For example, cuHALLaR is able to solve a matrix completion SDP instance with matrix variable of size 8 million, and approximately 300 million constraints, in just 142 seconds.
@@ 1.2 Notation
Let $\mathbb{R}^{n}$ be the space of $n$ dimensional vectors, $\mathbb{R}^{n\times r}$ the space of $n\times r$ matrices, and $\mathbb{S}^{n}$ the space of symmetric $n\times n$ matrices. Let $\mathbb{R}_{++}^{n}$ (resp., $\mathbb{R}^{n}_{+}$) denote the convex cone in $\mathbb{R}^{n}$ of vectors with positive (resp., nonnegative) entries and $\langle\cdot,\cdot\rangle$ (resp. $\|\cdot\|$) be the Euclidean inner product (resp. norm) on $\mathbb{R}^{n}$. Given matrices $A_{1},\ldots,A_{m}\in\mathbb{S}^{n}$, let $\mathcal{A}:\mathbb{S}^{n}\rightarrow\mathbb{R}^{m}$ denote the operator defined as $\mathcal{A}(X)]_{i}=A_{i}\bullet X$ for every $i=1,\ldots,m$. The adjoint operator $\mathcal{A}^{*}:\mathbb{R}^{m}\rightarrow\mathbb{S}^{n}$ of $\mathcal{A}$ is given by $\mathcal{A}^{*}(\lambda)=\sum_{i=1}^{m}\lambda_{i}A_{i}$. The minimum eigenvalue of a matrix $Q\in\mathbb{S}^{n}$ is denoted by $\lambda_{\min}(Q)$, and $v_{\min}(Q)$ denotes a corresponding eigenvector of unit norm.
@@ 2 Overview of HALLaR
As mentioned in the introduction, HALLaR is an inexact augmented Lagrangian method that approximately solves the pair of semidefinite programs (P) and (D). Given a tolerance pair $(\epsilon_{p},\epsilon_{pd})\in\mathbb{R}^{2}_{++}$, the method aims to find an $(\epsilon_{p},\epsilon_{pd})$-optimal solution, i.e., a triple $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}_{\tau}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$ that satisfies
$\|\mathcal{A}\bar{X}-b\|\leq\epsilon_{p},\quad|C\bullet\bar{X}+b^{\top}\bar{p}+\tau^{2}\bar{\theta}|\leq\epsilon_{pd},\quad C+\mathcal{A}^{*}\bar{p}+\bar{\theta}I\succeq 0,\quad\bar{\theta}\geq 0.$ (4)
When the tolerances are zero, these relations reduce to the standard optimality conditions for the primal-dual pair. Note that the final two inequalities in (4) enforce dual feasibility, a property that HALLaR maintains at every iteration. Being an inexact AL method, HALLaR generates sequences $\{X_{t}\}$ and $\{p_{t}\}$ according to the following updates
$X_{t}\quad$ $\approx\quad\operatorname*{arg\,min}_{X}\quad\{\mathcal{L}_{\beta_{t}}(X;p_{t-1})\;:\;X\in\Delta^{n}_{\tau}\},$ (5a)
@@
is the AL function. The key part of HALLaR is an efficient method, called the hybrid low-rank method (HLR), for finding a near-optimal solution $X_{t}$ of the AL subproblem (5a). HLR never forms $X_{t}$ but outputs a low-rank factor $U_{t}\in B_{\tau}^{s_{t}}$ such that $X_{t}=U_{t}U_{t}^{\top}$, where $B_{\tau}^{s_{t}}$ is as in (3) and hopefully $s_{t}\ll n$. The next paragraph briefly describes HLR in more detail for solving (5a).
Given a pair $(s,\tilde{Y})\in\mathbb{Z}_{+}\times B_{\tau}^{s}$, which is initially set to $(s,\tilde{Y})=(s_{t-1},U_{t-1})$, a general iteration of HLR performs the following steps: i) it computes a suitable stationary point $Y$ of the nonconvex reformulation of (5a) obtained through the change of variable $X=UU^{\top}$, namely:
$\min_{U}\quad\left\{g_{t}(U)=\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad\|U\|_{F}\leq\tau,\quad U\in\mathbb{R}^{n\times s}\right\};$ ($\mathrm{L}_{r}$)
[…]
$gt(Y~).R\in\nabla g_{t}(Y)+N_{B^{s}_{\tau}}(Y),\quad\|R\|_{F}\leq\rho$$\quad$ $\leq g_{t}(\tilde{Y})$. (7)
@@
$R\in 2[C+\mathcal{A}^{*}(p_{t-1}+\beta_{t}(\mathcal{A}(YY^{\top})-b))]Y+N_{B^{s}_{\tau}}(Y),\quad\|R\|_{F}\leq\rho.$
@@
Algorithm 1 HALLaR.
1:Input: Let initial points $(U_{0},p_{0})\in B^{s_{0}}_{\tau}\times\mathbb{R}^{m}$, tolerance pair $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})\in\mathbb{R}^{2}_{++}$, let $\{\beta_{t}\}_{t\geq 1}$ be a sequence of positive reals, and let $\left\{\epsilon_{t}\right\}_{t\geq 1}$ be a decreasing sequence of positive reals converging to 0.
2:Output: $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}_{\tau}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$, an $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})$-solution of the pair of SDPs (P) and (D).
3:set $t\leftarrow 1$;
4:set $\tilde{Y}=U_{t-1}$, $s=s_{t-1}$, $Y=0_{n\times s}$, $G=0_{n\times n}$, and $\theta=\infty$;
5:while $G\bullet(YY^{\top})+\tau^{2}\theta>\epsilon_{t}$ do $\triangleright$ (HLR method)
6:  call a nonconvex solver with initial point $\tilde{Y}$, tolerance $\epsilon_{t}$, and function $\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})+\delta_{B_{\tau}^{s}}(U)$ to find a point $Y\in B^{s}_{\tau}$ that is an $\epsilon_{t}$-approximate stationary solution (according to the criterion in (7)) of
$\min_{U}\quad\left\{\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad\|U\|_{F}\leq\tau,\quad U\in\mathbb{R}^{n\times s}\right\};$ (8)
7:  compute
$G:=\nabla\left[\mathcal{L}_{\beta_{t}}\left(\,\cdot\,;p_{t-1}\right)\right]\left(YY^{\top}\right)\in\mathbb{S}^{n}$ (9)
@@
8:  if $G\bullet(YY^{\top})+\tau^{2}\theta\leq\epsilon_{t}$ then
9:   break while loop and go to step 12 below;
10:  end if
11:  compute
$\alpha=\min\left\{\frac{G\bullet(YY^{\top})+\tau^{2}\theta}{\beta_{t}\|\mathcal{A}(YY^{\top})-\mathcal{A}(\tau^{2}yy^{\top})\|^{2}},\;1\right\}$ (11)
and set
@@
12:end while
13:set $(U_{t},\theta_{t},s_{t})=(Y,\theta,s)$;
14:set
$p_{t}=p_{t-1}+\beta_{t}(\mathcal{A}(U_{t}U_{t}^{\top})-b);$ (13)
15:if
$\|\mathcal{A}(U_{t}U_{t}^{\top})-b\|\leq\epsilon_{\mathrm{p}},\quad|C\bullet(U_{t}U_{t}^{\top})-(-b^{\top}p_{t}-\tau^{2}\theta_{t})|\leq\epsilon_{\mathrm{pd}}$ (14)
then return $(\bar{U},\bar{p},\bar{\theta})=(U_{t},p_{t},\theta_{t})$;
16:set $t=t+1$ and go to step 3.
Several remarks about each of the steps of HALLaR are now given. First, $t$ is the iteration count for HALLaR. Second, in practice we choose the sequence $\{\beta_{t}\}_{t\geq 1}$ in an adaptive manner based on some of the ideas of the LANCELOT method [17]. Third, the while loop in lines 5 to 12 of HALLaR corresponds to the HLR method, which is used by HALLaR to approximately solve its AL subproblems in (5a). In line 4, the triple $(Y,G,\theta)$ is set as $(0_{n\times s},0_{n\times n},\infty)$ to ensure that the HLR method is always executed during each iteration of HALLaR. Fourth, the nonconvex solver in line 6 of HALLaR will typically make several evaluations of $\nabla_{U}\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})$, which is easily seen to be equal to
$\nabla_{U}\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1}):=2\left[C+\mathcal{A}^{*}\left(p_{t-1}+\beta_{t}\left(\mathcal{A}(UU^{\top})-b\right)\right)\right]U.$ (15)
Fifth, it follows from the definition of $\mathcal{L}_{\beta_{t}}(\cdot;p_{t-1})$ that $G$ in (9) can be written as
$G:=C+\mathcal{A}^{*}\left(p_{t-1}+\beta_{t}\left(\mathcal{A}(YY^{\top})-b\right)\right).$ (16)
Sixth, after the HLR method approximately solves an AL subproblem, HALLaR updates the Lagrange multiplier in line 14 and checks for termination in line 15. Lastly, the termination condition on this line does not check for near dual feasibility since it can be shown that for every $t\geq 1$,
@@
and hence that HALLaR always generates feasible dual iterates $(p_{t},\theta_{t},S_{t})$.
@@
We now comment on the steps performed within the while loop (lines 5 to 12) in light of the three steps of the HALLaR outline given just before its formal description. Step 6 implements step i) of the outline. Steps 7 through 9 implement step ii). Finally, step 11 implements step iii).
We now give an interpretation of step 11 of HALLaR relative to the $X$-space. The Frank-Wolfe procedure generates a new iterate by moving from the current point $YY^{\top}$ towards an extreme point of the feasible set $\Delta^{n}_{\tau}$, which is of the form $\tau^{2}yy^{\top}$. The stepsize $\alpha$ computed in (11) corresponds to solving
$\operatorname*{arg\,min}_{\hat{\alpha}\in[0,1]}\left\{\mathcal{L}_{\beta_{t}}\left((1-\hat{\alpha})YY^{\top}+\hat{\alpha}(\tau yy^{\top});p_{t-1}\right)\right\}.$